What is the probability that a randomly chosen four letter word will begin with HA
Answer #1

We have 26 letters

n= Choosing 4 letters from 26 letters as 26c4 = 14950 ways

m= word started with ha

m = choosing ha has 1way of possiblility and the remaining two letters are choosen from 26c2 ways=325

Then p (4 letter word)= m/n = 325/14950 =0.0217